OpenStudy (anonymous):

what is the velum of a prism where the base is 49-49-90 triangle with a leg of 5 inches in length? the height of the prism is 1.8

OpenStudy (raden):

volume = base area * height

OpenStudy (raden):

base area = area of triangle = 1/2 * base * height of triangle = 1/2 * 5 * 5 = 25/2

OpenStudy (raden):

therefore, volume the prism = 25/2 * 1.8 = ....

OpenStudy (anonymous):

22.5

OpenStudy (raden):

yes, correct

OpenStudy (anonymous):

awh well thanks helping me more than my teacher (:

OpenStudy (raden):

welcome btw, that's should 45-45-90, right ?

OpenStudy (anonymous):

hmm yes it is sorry!!

OpenStudy (raden):

hehe, nope :)